Ask
your own question, for FREE!
Mathematics
7 Online
A box contains 4 black checkers and 8 red checkers. A checker is drawn and replaced. Then a second checker is drawn. Find the probability that both checkers are red.
Still Need Help?
Join the QuestionCove community and study together with friends!
multiply the probability that the first one is red, times the probability that the second one is red
okay
\[\frac{8}{12}\times \frac{8}{12}\] or maybe \[\frac{2}{3}\times \frac{2}{3}\]
for the first red the prob is 8/12 and for the second too it is 8/12 (note replacement done) hence th reqd probability is 8/12 *8/12=4/9
ya it's 4/9 that's what I got thanks guys
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
18 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal